reduce rational expression terms; there are no non-zeros

1) a square -b square over a-b
2) 3x-6 over x square- 4x+4

Answer by checkley77(12844)

1) a square -b square over a-b
(A^2-B^2)/(A-B)
(A+B)(A-B)/(A-B) CANCEL THE (A-B) TERMS.
A+B ANSWER.
2) 3x-6 over x square- 4x+4
3(X-2)/X^2-4X+4 The 4 is changed to the (4=-2*-2). The (x^2=x*x) & the (-4x=-2x-2x)
3(X-2)/(X-2)(X-2) CANCEL THE (X-2) TERMS.
3/(X-2) ANSWER.